# Orbifold modifications of complex analytic varieties

### v4

@@ 6.B. Construction of covering spaces
We notice that $\widehat{S}\cong S$. From the construction of $W$ in Lemma 6.10, we see that if $\widehat{H}$ is the preimage of $H$ in $\widehat{S}$, then the data $W$, $\rho^{-1}(\Delta)|_{W}$, $\widehat{S}$, $\widehat{H}$ and $G$ satisfy the items (1) and (2) of Setup 6.6. If $o_{\widehat{S}}$ is the preimage of $o_{S}$ in $\widehat{S}$, then there is an open neighborhood $S_{1}$ of $o_{\widehat{S}}$ in $\widehat{S}$, which plays the same role for $o_{\widehat{S}}$ as $S$ for $o_{S}$ in the item (4) of Setup 6.6. By applying Lemma 6.9 to $W$, there is an open embedding $\hat{\varphi}\colon U\times S_{1}^{\circ}\to W$ whose restriction on $\{\mathbf{0}_{3}\}\times S_{1}^{\circ}$ is the identity map to $S_{1}^{\circ}$, where $S_{1}^{\circ}=S_{1}\setminus\widehat{H}$. Since there are natural isomorphisms
@@
We recall that $\varphi\colon U\times S^{\circ}\to V$ is an isomorphism. Since there is a natural isomorphism $\pi_{1}(U_{\operatorname{\textsubscript{\rm sm}}})\cong\pi_{1}(Y_{\operatorname{\textsubscript{\rm sm}}})$, we deduce that
$\varphi^{-1}_{*}\colon\pi_{1}(V_{\operatorname{\textsubscript{\rm sm}}})\to\pi_{1}(Y_{\operatorname{\textsubscript{\rm sm}}}\times\check{S}^{\circ})$
is an isomorphism, and we obtain that $\rho_{*}\colon\pi_{1}(W^{\circ}_{\operatorname{\textsubscript{\rm sm}}})\to\pi_{1}(V_{\operatorname{\textsubscript{\rm sm}}})$ is surjective. Moreover, the kernel of $\iota$ is equal to the one of $\rho_{*}$.
Assume that we have a morphism $\psi\colon X^{\circ}\to Y\times\check{S}^{\circ}$ as in the second part of the lemma. Then the morphism
$\psi\circ\varphi\colon U\times S^{\circ}\to Y\times\check{S}^{\circ}$
@@
$\psi_{*}\colon\pi_{1}(V_{\operatorname{\textsubscript{\rm sm}}})\to\pi_{1}(Y_{\operatorname{\textsubscript{\rm sm}}}\times\check{S}^{\circ})$
is an isomorphism. Since $\rho_{*}\colon\pi_{1}(W^{\circ}_{\operatorname{\textsubscript{\rm sm}}})\to\pi_{1}(V_{\operatorname{\textsubscript{\rm sm}}})$ is surjective, we deduce that $(\psi\circ\rho)_{*}$ is surjective. Furthermore, the kernel of $(\psi\circ\rho)_{*}$ is the one of $\rho_{*}$, which is also the kernel of $\iota$. This completes the proof of the lemma. ∎
@@
With the notation in Lemma 6.11, let $\hat{o}\in W\setminus\widehat{X}^{\circ}_{\operatorname{\textsubscript{\rm sm}}}$ be a point and let $o=\rho(\hat{o})\in X$. Let $\Sigma\subseteq X$ be an open neighborhood of $o$, and let $\Sigma^{\circ}=\Sigma\cap X^{\circ}$. Assume that for $i=1,2$, there is an open embedding $\psi_{i}\colon\Sigma^{\circ}\to Y\times(\mathbb{C}^{n}\setminus H^{i})$ satisfying the following properties. Each $H^{i}$ is the union of some coordinate hyperplanes in $\mathbb{C}^{n}$, the restriction $\psi_{i}|_{\check{S}\cap\Sigma^{\circ}}$ has images contained in $\{\mathbf{0}_{3}\}\times\mathbb{C}^{n}$, and it extends to an open embedding from $\check{S}\cap\Sigma$ to $\mathbb{C}^{n}$ such that $\psi_{i}(\check{S}\cap\Sigma^{\circ})=\psi_{i}(\check{S}\cap\Sigma)\setminus H^{i}$.
@@
We first assume that the point $\hat{o}$ is in $W^{\circ}$. Then $\hat{o}\in W^{\circ}_{\operatorname{\textsubscript{\rm sing}}}$. Since $\rho$ is an isomorphism on $W^{\circ}$, we deduce that $o\in S^{\circ}\subseteq X^{\circ}$. It follows that $k^{\prime}=0$. We note that $X$ has the same type of singularities at points of $S^{\circ}$. Hence there is a contractible open neighborhood $S^{\prime}$ of $o$ in $S^{\circ}$, an open neighborhood $V^{\prime}$ of $o$ in $\Sigma^{\circ}$, and an open neighborhood $Y^{\prime}$ of $\mathbf{0}_{3}$ in $Y$, such that $V^{\prime}\cong Y^{\prime}\times S^{\prime}$. We may assume further that the natural morphism $\pi_{1}(Y^{\prime}_{\operatorname{\textsubscript{\rm sm}}})\to\pi_{1}(Y_{\operatorname{\textsubscript{\rm sm}}})$ is isomorphic. Let $\widehat{\Sigma}$ be the preimage of $V^{\prime}$ in $\widehat{X}$. Then we have $\hat{o}\in\widehat{\Sigma}\subseteq\widehat{X}^{\circ}$ and $\rho$ induces an isomorphism from $\widehat{\Sigma}$ to $V^{\prime}$. In particular, $\rho(\widehat{\Sigma})\cap\check{S}^{\circ}=S^{\prime}$, and there are natural isomorphisms
@@
It remains to treat the case when $\hat{o}\in W\setminus W^{\circ}=W\cap\rho^{-1}(\Delta)$. This assumption implies that $o\in H\cap S$ by Lemma 6.10. We recall that $\theta\colon M\to\check{S}$ is the projection sending a point $(x,y,z,s)\in M$ to $s\in\check{S}$. Let $\hat{\theta}\colon\widehat{M}\to\check{S}$ be the composition $\theta\circ\rho$. By Lemma 6.9, there is an open neighborhood $S^{\prime}$ of $o$ in $S$, such that
@@
$\pi_{1}(V^{\prime}_{\operatorname{\textsubscript{\rm sm}}})\to\pi_{1}(Y_{\operatorname{\textsubscript{\rm sm}}}\times(\mathbb{C}^{n}\setminus H^{i}))$
induced by $\psi_{i}$ is injective. We then deduce that the kernel of the morphism from $\pi_{1}(\widehat{\Sigma}_{\operatorname{\textsubscript{\rm sm}}}\cap\widehat{X}^{\circ})$ to $\pi_{1}(Y_{\operatorname{\textsubscript{\rm sm}}}\times(\mathbb{C}^{n}\setminus H^{i}))$ induced by $\psi_{i}\circ\rho$ is the same as the one of $\rho_{*}\colon\pi_{1}(\widehat{\Sigma}_{\operatorname{\textsubscript{\rm sm}}}\cap\widehat{X}^{\circ})\to\pi_{1}(V^{\prime}_{\operatorname{\textsubscript{\rm sm}}})$, which is independent of $i$. Hence we obtain the item (1).
Lemma 6.4 also implies that the image of $(\psi_{i})_{*}(\pi_{1}(V^{\prime}_{\operatorname{\textsubscript{\rm sm}}}))$ is the same as the image of $\pi_{1}(Y_{\operatorname{\textsubscript{\rm sm}}}\times\psi_{i}(S^{\prime}\cap\check{S}^{\circ}))$ induced by the natural inclusion. Furthermore, by the same argument as in the first two paragraph of the proof of Lemma 6.11, we see that the morphism $\rho_{*}\colon\pi_{1}(\widehat{\Sigma}_{\operatorname{\textsubscript{\rm sm}}}\cap\widehat{X}^{\circ})\to\pi_{1}(V^{\prime}_{\operatorname{\textsubscript{\rm sm}}})$ is surjective. Hence the item (2) holds as well. This completes the proof of the lemma. ∎
@@
$\iota=(\varphi^{-1}\circ\rho)_{*}\colon\pi_{1}(W^{\circ}_{\operatorname{\textsubscript{\rm sm}}})\to\pi_{1}(Y_{\operatorname{\textsubscript{\rm sm}}}\times\check{S}^{\circ}).$
We note that $\pi_{1}(Y_{\operatorname{\textsubscript{\rm sm}}}\times\check{S}^{\circ})\cong G_{Y}\times\mathbb{Z}^{k}$, where $G_{Y}$ is the fundamental group of $Y_{\operatorname{\textsubscript{\rm sm}}}$, which is finite. Let $\mathcal{N}\subseteq\pi_{1}(Y_{\operatorname{\textsubscript{\rm sm}}}\times\check{S}^{\circ})$ be the normal subgroup defined in Lemma 6.5, and let $\mathcal{H}=\iota^{-1}(\mathcal{N})$. Then $\mathcal{H}$ is a normal subgroup of $\pi_{1}(W^{\circ}_{\operatorname{\textsubscript{\rm sm}}})$, and it determines a finite Galois étale cover of $W^{\circ}_{\operatorname{\textsubscript{\rm sm}}}$. Let $\nu\colon\overline{W}\to W$ be the induced finite cover with $\overline{W}$ normal, see Theorem 2.4. Then $\nu$ is Galois and quasi-étale over $W^{\circ}$. We have hence proved the items (1) and (2).
We will show that $\overline{W}^{\circ}:=\nu^{-1}(W^{\circ})$ is smooth. Let $Z\to Y\times\check{S}^{\circ}$ be the finite quasi-étale morphism induced by the subgroup $\mathcal{N}\subseteq\pi_{1}(Y_{\operatorname{\textsubscript{\rm sm}}}\times\check{S}^{\circ})$, see Theorem 2.4. Then it factors through $\overline{Y}\times\check{S}^{\circ}\to Y\times\check{S}^{\circ}$, where $\overline{Y}\to Y$ is the quasi-étale cover induced by the universal cover of $Y_{\operatorname{\textsubscript{\rm sm}}}$. In particular, $\overline{Y}$ is smooth. The Zariski’s purity theorem then implies that $Z$ is smooth. We note that $\nu|_{\overline{W}^{\circ}}$ is the basechange of $\varphi^{-1}\circ\rho\colon W^{\circ}\to Y\times\check{S}^{\circ}$ over $Z\to Y\times\check{S}^{\circ}$.
@@
To prove the item (4), it is enough to show that $\mathcal{H}$ is $G$-invariant. Let $g\in G$ be an element. Then $g^{-1}(\mathcal{H})$ is the preiamge of $\mathcal{H}$ under the morphism of fundamental groups induced by $g\colon W\to W$. Let
$\iota^{\prime}:=(g^{-1}\circ\varphi^{-1}\circ\rho\circ g)_{*}\colon\pi_{1}(W^{\circ}_{\operatorname{\textsubscript{\rm sm}}})\to\pi_{1}(Y_{\operatorname{\textsubscript{\rm sm}}}\times\check{S}^{\circ}),$
@@
$\psi=g^{-1}\circ\varphi^{-1}\circ g\colon X^{\circ}\to Y\times\check{S}^{\circ},$
@@
Next we assume that $\hat{o}$ is not contained in $\widehat{X}^{\circ}_{\operatorname{\textsubscript{\rm sm}}}$. Then $o=\rho(\hat{o})\in S$, see Lemma 6.10. As explained in the paragraph after Proposition 6.7, by assumption, there is an open neighborhood $X^{\prime}$ of $o=\rho(\hat{o})$ containing $\rho(W^{\prime})$, there is an open embedding $\psi\colon(X^{\prime}\cap X^{\circ})\to Y\times(\mathbb{C}^{n}\setminus H^{\prime})$, where $H^{\prime}$ is the union of some coordinate hyperplanes in $\mathbb{C}^{n}$, such that the following properties hold. The restriction of $\psi$ on $X^{\prime}\cap\check{S}^{\circ}$ has images contained in $\{\mathbf{0}_{3}\}\times\mathbb{C}^{n}$. It extends to an open embedding from $X^{\prime}\cap\check{S}$ to $\mathbb{C}^{n}$, such that $\psi(X^{\prime}\cap\check{S}^{\circ})=\psi(X^{\prime}\cap\check{S})\setminus H^{\prime}$. Furthermore, $\nu^{\prime}$ is induced by the kernel of the composition of the following (conjugacy classes of) morphisms of groups
@@
$\sigma_{1}=(\varphi^{-1}\circ\rho)_{*}\colon\pi_{1}((W^{\prime\prime}\cap\widehat{X}^{\circ})_{\operatorname{\textsubscript{\rm sm}}})\to\pi_{1}(Y_{\operatorname{\textsubscript{\rm sm}}}\times\check{S}^{\circ})$
@@
$\sigma_{2}=(\psi\circ\rho)_{*}\colon\pi_{1}((W^{\prime\prime}\cap\widehat{X}^{\circ})_{\operatorname{\textsubscript{\rm sm}}})\to\pi_{1}(Y_{\operatorname{\textsubscript{\rm sm}}}\times(\mathbb{C}^{n}\setminus H^{\prime}))$
have the same kernel. Their images are both normal subgroups and both isomorphic to $G_{Y}\times\mathbb{Z}^{k^{\prime\prime}}$, where $k^{\prime\prime}$ is the number of the components of $H$ which pass through $o$. Hence there is a surjective morphism $\mu\colon\pi_{1}((W^{\prime\prime}\cap\widehat{X}^{\circ})_{\operatorname{\textsubscript{\rm sm}}})\to G_{Y}\times\mathbb{Z}^{k^{\prime\prime}}$, such that $\sigma_{i}=\chi_{i}\circ\mu$ for $i=1,2$, where
$\chi_{1}\colon G_{Y}\times\mathbb{Z}^{k^{\prime\prime}}\to\pi_{1}(Y_{\operatorname{\textsubscript{\rm sm}}}\times\check{S}^{\circ})$
@@
$\chi_{2}\colon G_{Y}\times\mathbb{Z}^{k^{\prime\prime}}\to\pi_{1}(Y_{\operatorname{\textsubscript{\rm sm}}}\times(\mathbb{C}^{n}\setminus H^{\prime}))$
@@
We notice that, if ${\nu^{-1}(W^{\prime\prime})}_{\circ}$ is a connected component of ${\nu^{-1}(W^{\prime\prime})}$, then $\nu|_{\nu^{-1}(W^{\prime\prime})_{\circ}}\colon{\nu^{-1}(W^{\prime\prime})}_{\circ}\to W^{\prime\prime}$ is induced by the kernel of the composition of the following sequence of morphisms of groups
@@
which is equal to $\mu^{-1}(\chi_{1}^{-1}(\mathcal{N}))$. Since the cokernel of $\chi_{1}$ is isomorphic to $\mathbb{Z}^{k-k^{\prime}}$, the subgroup $\mathcal{N}^{\prime\prime}:=\chi_{1}^{-1}(\mathcal{N})$ of $G_{Y}\times\mathbb{Z}^{k^{\prime\prime}}$ is the canonical subgroup defined in Lemma 6.5. Then $\nu|_{\nu^{-1}(W^{\prime\prime})_{\circ}}$ is induced by the subgroup $\mu^{-1}(\mathcal{N^{\prime\prime}})$. Similarly, if ${(\nu^{\prime})^{-1}(W^{\prime\prime})}_{\circ}$ is a connected component of ${(\nu^{\prime})^{-1}(W^{\prime\prime})}$, then the finite morphism $\nu^{\prime}|_{(\nu^{\prime})^{-1}(W^{\prime\prime})_{\circ}}\colon{(\nu^{\prime})^{-1}(W^{\prime\prime})}_{\circ}\to W^{\prime\prime}$ is induced by the subgroup $\mu^{-1}(\chi_{2}^{-1}(\mathcal{N^{\prime}}))$. The same argument implies that $\chi_{2}^{-1}(\mathcal{N^{\prime}})$ is the canonical subgroup of $G_{Y}\times\mathbb{Z}^{k^{\prime\prime}}$ in Lemma 6.5. Hence $\chi_{2}^{-1}(\mathcal{N})=\mathcal{N}^{\prime\prime}$, and we deduce that $\nu|_{\nu^{-1}(W^{\prime\prime})_{\circ}}$ is isomorphic to $\nu^{\prime}|_{(\nu^{\prime})^{-1}(W^{\prime\prime})_{\circ}}$. This completes the proof of the proposition. ∎
@@
[…]
@@
where $\overline{s}\in\overline{S}$ is a point and $p\colon\overline{S}\to\check{S}$ is the natural projection. From the construction, we deduce directly that the vanishing orders of $\alpha\circ p$ and $\beta\circ p$ along each $\overline{H}_{i}$ are divisible by 6. This completes the proof of the lemma. ∎

### v6

@@ 6.B. Construction of covering spaces
We notice that $\widehat{S}\cong S$. From the construction of $W$ in Lemma 6.10, we see that if $\widehat{H}$ is the preimage of $H$ in $\widehat{S}$, then the data $W$, $\rho^{-1}(\Delta)|_{W}$, $\widehat{S}$, $\widehat{H}$ and $G$ satisfy the items (1) and (2) of Setup 6.6. If $o_{\widehat{S}}$ is the preimage of $o_{S}$ in $\widehat{S}$, then there is an open neighborhood $S_{1}$ of $o_{\widehat{S}}$ in $\widehat{S}$, which plays the same role for $o_{\widehat{S}}$ as $S$ for $o_{S}$ in the item (4) of Setup 6.6. By applying Lemma 6.9 to $W$, there is an open embedding $\hat{\varphi}\colon U\times S_{1}^{\circ}\rightarrow W$ whose restriction on $\{\mathbf{0}_{3}\}\times S_{1}^{\circ}$ is the identity map to $S_{1}^{\circ}$, where $S_{1}^{\circ}=S_{1}\setminus\widehat{H}$. Since there are natural isomorphisms
@@
We recall that $\varphi\colon U\times S^{\circ}\rightarrow V$ is an isomorphism. Since there is a natural isomorphism $\pi_{1}(U_{\operatorname{\textsubscript{\rm sm}}})\cong\pi_{1}(Y_{\operatorname{\textsubscript{\rm sm}}})$, we deduce that
$\varphi^{-1}_{*}\colon\pi_{1}(V_{\operatorname{\textsubscript{\rm sm}}})\rightarrow\pi_{1}(Y_{\operatorname{\textsubscript{\rm sm}}}\times\check{S}^{\circ})$
is an isomorphism, and we obtain that $\rho_{*}\colon\pi_{1}(W^{\circ}_{\operatorname{\textsubscript{\rm sm}}})\rightarrow\pi_{1}(V_{\operatorname{\textsubscript{\rm sm}}})$ is surjective. Moreover, the kernel of $\iota$ is equal to the one of $\rho_{*}$.
Assume that we have a morphism $\psi\colon X^{\circ}\rightarrow Y\times\check{S}^{\circ}$ as in the second part of the lemma. Then the morphism
$\psi\circ\varphi\colon U\times S^{\circ}\rightarrow Y\times\check{S}^{\circ}$
@@
$\psi_{*}\colon\pi_{1}(V_{\operatorname{\textsubscript{\rm sm}}})\rightarrow\pi_{1}(Y_{\operatorname{\textsubscript{\rm sm}}}\times\check{S}^{\circ})$
is an isomorphism. Since $\rho_{*}\colon\pi_{1}(W^{\circ}_{\operatorname{\textsubscript{\rm sm}}})\rightarrow\pi_{1}(V_{\operatorname{\textsubscript{\rm sm}}})$ is surjective, we deduce that $(\psi\circ\rho)_{*}$ is surjective. Furthermore, the kernel of $(\psi\circ\rho)_{*}$ is the one of $\rho_{*}$, which is also the kernel of $\iota$. This completes the proof of the lemma. ∎
@@
With the notation in Lemma 6.11, let $\hat{o}\in W\setminus\widehat{X}^{\circ}_{\operatorname{\textsubscript{\rm sm}}}$ be a point and let $o=\rho(\hat{o})\in X$. Let $\Sigma\subseteq X$ be an open neighborhood of $o$, and let $\Sigma^{\circ}=\Sigma\cap X^{\circ}$. Assume that for $i=1,2$, there is an open embedding $\psi_{i}\colon\Sigma^{\circ}\rightarrow Y\times(\mathbb{C}^{n}\setminus H^{i})$ satisfying the following properties. Each $H^{i}$ is the union of some coordinate hyperplanes in $\mathbb{C}^{n}$, the restriction $\psi_{i}|_{\check{S}\cap\Sigma^{\circ}}$ has images contained in $\{\mathbf{0}_{3}\}\times\mathbb{C}^{n}$, and it extends to an open embedding from $\check{S}\cap\Sigma$ to $\mathbb{C}^{n}$ such that $\psi_{i}(\check{S}\cap\Sigma^{\circ})=\psi_{i}(\check{S}\cap\Sigma)\setminus H^{i}$.
@@
We first assume that the point $\hat{o}$ is in $W^{\circ}$. Then $\hat{o}\in W^{\circ}_{\operatorname{\textsubscript{\rm sing}}}$. Since $\rho$ is an isomorphism on $W^{\circ}$, we deduce that $o\in S^{\circ}\subseteq X^{\circ}$. It follows that $k^{\prime}=0$. We note that $X$ has the same type of singularities at points of $S^{\circ}$. Hence there is a contractible open neighborhood $S^{\prime}$ of $o$ in $S^{\circ}$, an open neighborhood $V^{\prime}$ of $o$ in $\Sigma^{\circ}$, and an open neighborhood $Y^{\prime}$ of $\mathbf{0}_{3}$ in $Y$, such that $V^{\prime}\cong Y^{\prime}\times S^{\prime}$. We may assume further that the natural morphism $\pi_{1}(Y^{\prime}_{\operatorname{\textsubscript{\rm sm}}})\rightarrow\pi_{1}(Y_{\operatorname{\textsubscript{\rm sm}}})$ is isomorphic. Let $\widehat{\Sigma}$ be the preimage of $V^{\prime}$ in $\widehat{X}$. Then we have $\hat{o}\in\widehat{\Sigma}\subseteq\widehat{X}^{\circ}$ and $\rho$ induces an isomorphism from $\widehat{\Sigma}$ to $V^{\prime}$. In particular, $\rho(\widehat{\Sigma})\cap\check{S}^{\circ}=S^{\prime}$, and there are natural isomorphisms
@@
It remains to treat the case when $\hat{o}\in W\setminus W^{\circ}=W\cap\rho^{-1}(\Delta)$. This assumption implies that $o\in H\cap S$ by Lemma 6.10. We recall that $\theta\colon M\rightarrow\check{S}$ is the projection sending a point $(x,y,z,s)\in M$ to $s\in\check{S}$. Let $\hat{\theta}\colon\widehat{M}\rightarrow\check{S}$ be the composition $\theta\circ\rho$. By Lemma 6.9, there is an open neighborhood $S^{\prime}$ of $o$ in $S$, such that
@@
$\pi_{1}(V^{\prime}_{\operatorname{\textsubscript{\rm sm}}})\rightarrow\pi_{1}(Y_{\operatorname{\textsubscript{\rm sm}}}\times(\mathbb{C}^{n}\setminus H^{i}))$
induced by $\psi_{i}$ is injective. We then deduce that the kernel of the morphism from $\pi_{1}(\widehat{\Sigma}_{\operatorname{\textsubscript{\rm sm}}}\cap\widehat{X}^{\circ})$ to $\pi_{1}(Y_{\operatorname{\textsubscript{\rm sm}}}\times(\mathbb{C}^{n}\setminus H^{i}))$ induced by $\psi_{i}\circ\rho$ is the same as the one of $\rho_{*}\colon\pi_{1}(\widehat{\Sigma}_{\operatorname{\textsubscript{\rm sm}}}\cap\widehat{X}^{\circ})\rightarrow\pi_{1}(V^{\prime}_{\operatorname{\textsubscript{\rm sm}}})$, which is independent of $i$. Hence we obtain the item (1).
Lemma 6.4 also implies that the image of $(\psi_{i})_{*}(\pi_{1}(V^{\prime}_{\operatorname{\textsubscript{\rm sm}}}))$ is the same as the image of $\pi_{1}(Y_{\operatorname{\textsubscript{\rm sm}}}\times\psi_{i}(S^{\prime}\cap\check{S}^{\circ}))$ induced by the natural inclusion. Furthermore, by the same argument as in the first two paragraph of the proof of Lemma 6.11, we see that the morphism $\rho_{*}\colon\pi_{1}(\widehat{\Sigma}_{\operatorname{\textsubscript{\rm sm}}}\cap\widehat{X}^{\circ})\rightarrow\pi_{1}(V^{\prime}_{\operatorname{\textsubscript{\rm sm}}})$ is surjective. Hence the item (2) holds as well. This completes the proof of the lemma. ∎
@@
$\iota=(\varphi^{-1}\circ\rho)_{*}\colon\pi_{1}(W^{\circ}_{\operatorname{\textsubscript{\rm sm}}})\rightarrow\pi_{1}(Y_{\operatorname{\textsubscript{\rm sm}}}\times\check{S}^{\circ}).$
We note that $\pi_{1}(Y_{\operatorname{\textsubscript{\rm sm}}}\times\check{S}^{\circ})\cong G_{Y}\times\mathbb{Z}^{k}$, where $G_{Y}$ is the fundamental group of $Y_{\operatorname{\textsubscript{\rm sm}}}$, which is finite. Let $\mathcal{N}\subseteq\pi_{1}(Y_{\operatorname{\textsubscript{\rm sm}}}\times\check{S}^{\circ})$ be the normal subgroup defined in Lemma 6.5, and let $\mathcal{H}=\iota^{-1}(\mathcal{N})$. Then $\mathcal{H}$ is a normal subgroup of $\pi_{1}(W^{\circ}_{\operatorname{\textsubscript{\rm sm}}})$, and it determines a finite Galois étale cover of $W^{\circ}_{\operatorname{\textsubscript{\rm sm}}}$. Let $\nu\colon\overline{W}\rightarrow W$ be the induced finite cover with $\overline{W}$ normal, see Theorem 2.4. Then $\nu$ is Galois and quasi-étale over $W^{\circ}$. We have hence proved the items (1) and (2).
We will show that $\overline{W}^{\circ}:=\nu^{-1}(W^{\circ})$ is smooth. Let $Z\rightarrow Y\times\check{S}^{\circ}$ be the finite quasi-étale morphism induced by the subgroup $\mathcal{N}\subseteq\pi_{1}(Y_{\operatorname{\textsubscript{\rm sm}}}\times\check{S}^{\circ})$, see Theorem 2.4. Then it factors through $\overline{Y}\times\check{S}^{\circ}\rightarrow Y\times\check{S}^{\circ}$, where $\overline{Y}\rightarrow Y$ is the quasi-étale cover induced by the universal cover of $Y_{\operatorname{\textsubscript{\rm sm}}}$. In particular, $\overline{Y}$ is smooth. The Zariski’s purity theorem then implies that $Z$ is smooth. We note that $\nu|_{\overline{W}^{\circ}}$ is the basechange of $\varphi^{-1}\circ\rho\colon W^{\circ}\rightarrow Y\times\check{S}^{\circ}$ over $Z\rightarrow Y\times\check{S}^{\circ}$.
@@
To prove the item (4), it is enough to show that $\mathcal{H}$ is $G$-invariant. Let $g\in G$ be an element. Then $g^{-1}(\mathcal{H})$ is the preiamge of $\mathcal{H}$ under the morphism of fundamental groups induced by $g\colon W\rightarrow W$. Let
$\iota^{\prime}:=(g^{-1}\circ\varphi^{-1}\circ\rho\circ g)_{*}\colon\pi_{1}(W^{\circ}_{\operatorname{\textsubscript{\rm sm}}})\rightarrow\pi_{1}(Y_{\operatorname{\textsubscript{\rm sm}}}\times\check{S}^{\circ}),$
@@
$\psi=g^{-1}\circ\varphi^{-1}\circ g\colon X^{\circ}\rightarrow Y\times\check{S}^{\circ},$
@@
[…]
@@
$\sigma_{1}=(\varphi^{-1}\circ\rho)_{*}\colon\pi_{1}((W^{\prime\prime}\cap\widehat{X}^{\circ})_{\operatorname{\textsubscript{\rm sm}}})\rightarrow\pi_{1}(Y_{\operatorname{\textsubscript{\rm sm}}}\times\check{S}^{\circ})$
@@
$\sigma_{2}=(\psi\circ\rho)_{*}\colon\pi_{1}((W^{\prime\prime}\cap\widehat{X}^{\circ})_{\operatorname{\textsubscript{\rm sm}}})\rightarrow\pi_{1}(Y_{\operatorname{\textsubscript{\rm sm}}}\times(\mathbb{C}^{n}\setminus H^{\prime}))$
have the same kernel. Their images are both normal subgroups and both isomorphic to $G_{Y}\times\mathbb{Z}^{k^{\prime\prime}}$, where $k^{\prime\prime}$ is the number of the components of $H$ which pass through $o$. Hence there is a surjective morphism $\mu\colon\pi_{1}((W^{\prime\prime}\cap\widehat{X}^{\circ})_{\operatorname{\textsubscript{\rm sm}}})\rightarrow G_{Y}\times\mathbb{Z}^{k^{\prime\prime}}$, such that $\sigma_{i}=\chi_{i}\circ\mu$ for $i=1,2$, where
$\chi_{1}\colon G_{Y}\times\mathbb{Z}^{k^{\prime\prime}}\rightarrow\pi_{1}(Y_{\operatorname{\textsubscript{\rm sm}}}\times\check{S}^{\circ})$
@@
$\chi_{2}\colon G_{Y}\times\mathbb{Z}^{k^{\prime\prime}}\rightarrow\pi_{1}(Y_{\operatorname{\textsubscript{\rm sm}}}\times(\mathbb{C}^{n}\setminus H^{\prime}))$
@@
We notice that, if $\nu^{-1}(W^{\prime\prime})$$_{\circ}$ is a connected component of $\nu^{-1}(W^{\prime\prime})$}, then $\nu|_{\nu^{-1}(W^{\prime\prime})_{\circ}}\colon{\nu^{-1}(W^{\prime\prime})}_{\circ}\rightarrow W^{\prime\prime}$ is induced by the kernel of the composition of the following sequence of morphisms of groups
@@
which is equal to $\mu^{-1}(\chi_{1}^{-1}(\mathcal{N}))$. Since the cokernel of $\chi_{1}$ is isomorphic to $\mathbb{Z}^{k-k^{\prime}}$, the subgroup $\mathcal{N}^{\prime\prime}:=\chi_{1}^{-1}(\mathcal{N})$ of $G_{Y}\times\mathbb{Z}^{k^{\prime\prime}}$ is the canonical subgroup defined in Lemma 6.5. Then $\nu|_{\nu^{-1}(W^{\prime\prime})_{\circ}}$ is induced by the subgroup $\mu^{-1}(\mathcal{N^{\prime\prime}})$. Similarly, if $(\nu^{\prime})^{-1}(W^{\prime\prime})$$_{\circ}$ is a connected component of $(\nu^{\prime})^{-1}(W^{\prime\prime})$}, then the finite morphism $\nu^{\prime}|_{(\nu^{\prime})^{-1}(W^{\prime\prime})_{\circ}}\colon{(\nu^{\prime})^{-1}(W^{\prime\prime})}_{\circ}\rightarrow W^{\prime\prime}$ is induced by the subgroup $\mu^{-1}(\chi_{2}^{-1}(\mathcal{N^{\prime}}))$. The same argument implies that $\chi_{2}^{-1}(\mathcal{N^{\prime}})$ is the canonical subgroup of $G_{Y}\times\mathbb{Z}^{k^{\prime\prime}}$ in Lemma 6.5. Hence $\chi_{2}^{-1}(\mathcal{N})=\mathcal{N}^{\prime\prime}$, and we deduce that $\nu|_{\nu^{-1}(W^{\prime\prime})_{\circ}}$ is isomorphic to $\nu^{\prime}|_{(\nu^{\prime})^{-1}(W^{\prime\prime})_{\circ}}$. This completes the proof of the proposition. ∎
@@
With the notation in the item (1) of Setup 6.6, we let $\overline{S}=\check{S}[\sqrt[6]{T_{1}},...,\sqrt[6]{T_{n}}]$ and let $\overline{M}=M\times_{\check{S}}\overline{S}.$ We denote by $\overline{H}_{i}$ the preimage of $H_{i}$ in $\overline{S}$. Then $(\overline{S}_{i},\sum_{i=1}^{n}\overline{H_{i}})$ is snc and $M$ is a hypersurface in $\mathbb{D}^{3}\times\overline{S}$ defined by a function of the shape of Proposition 5.2 with respect to $(\overline{S}_{i},\sum_{i=1}^{n}\overline{H_{i}})$, such that the vanishing orders of the functions $\alpha$ and $\beta$ on $\overline{S}$ along each $\overline{H}_{i}$ are divisible by 6.
@@
where $\overline{s}\in\overline{S}$ is a point and $p\colon\overline{S}\rightarrow\check{S}$ is the natural projection. From the construction, we deduce directly that the vanishing orders of $\alpha\circ p$ and $\beta\circ p$ along each $\overline{H}_{i}$ are divisible by 6. This completes the proof of the lemma. ∎